OpenStudy (anonymous):

There are 6 brooms and 4 mops in a janitor's closet. What is the ratio of the number of brooms to the number of mops? A. 3/2 B. 3/5 C. 2/3 D. 5/3

OpenStudy (texaschic101):

6/4 = 3/2

OpenStudy (anonymous):

hmm i sure did that and got 1/2

OpenStudy (anonymous):

nvm i got it thanks

OpenStudy (texaschic101):

you have to pay attention to the wording....brooms to mops = 6/4 = 3/2. If it had said mops to brooms, it would have been 4/6 = 2/3